OpenStudy (anonymous):

Solve -4(x + 1) - 3 = -3(x - 4).

OpenStudy (anonymous):

your gonna have to multiply -4 by x and 1 so the answer to that would be -4x+1 then you do the same to the other side 3x-12 so that would be -4x+1 = 3x-12

OpenStudy (anonymous):

do you know what to do after that

OpenStudy (anonymous):

combine like terms?

OpenStudy (anonymous):

yeah but in this case theres nothing to combine

OpenStudy (anonymous):

iverse

OpenStudy (anonymous):

subtract the 3x by the -4x?

OpenStudy (anonymous):

yes subtract 3x from both sides... what would that be

OpenStudy (anonymous):

-7x

OpenStudy (anonymous):

so -7x+1 = -12 whats next

OpenStudy (anonymous):

wait wouldnt it be -4x+1+-3x+12?

OpenStudy (anonymous):

-4x+1=-3x+12

OpenStudy (anonymous):

no because your supposed to subtract 3x on both sides so 3x (on the right side) cancels out

OpenStudy (anonymous):

but in step 1 arent thee signs suppose to change when you distribute?

OpenStudy (anonymous):

these are the steps 1) distribute 2) combine like terms 3) move variables to once side ( witch we just did canceling the 3x) 4) subtract/addition thing lol i forgot that one 5) divide if needed to slove

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry that took a while those white boards are hard to use

OpenStudy (anonymous):

the 4 and 3 are suppose to be negative in the first one

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

opps i forgot to put the negative sign for 12

OpenStudy (anonymous):

but yeah do you knw what your supposed to do after that?

OpenStudy (anonymous):

im all confused hold on lol

OpenStudy (anonymous):

Solve -4(x + 1) - 3 = -3(x - 4). -4x-4-3 = -3x+12 -4x-7 = -3x+12 -x-7=12 -x=19 x=-19?

OpenStudy (anonymous):

ty for the help

OpenStudy (anonymous):

and medal

OpenStudy (anonymous):

np